Find three consecutive even integers such that the product of the first and second is four less than the square of the third.
Answer by stanbon(57203)

Find three consecutive even integers such that the product of the first and second is four less than the square of the third.
---
1st: 2x-2
2nd: 2x
3rd: 2x+2
-------------------
Equation:
(2x-2)(2x) = (2x+2)^2-4
---
4x^2-4x = 4x^2 + 8x + 4 -4
----
12x = 0
x = 0
---
1st: 2*0-2 = -2
2nd: 2x = 0
3rd: 2x+2 = 2
===================
Cheers,
Stan H.
===================